OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2 . y =1/2 x + 5 y = 1/2x - 3 y =1/2 x - 5 y = 1/2x + 3

OpenStudy (anonymous):

Use point slope form. y-y1=m(x-x1) y+4=1/2(x-2) y+4=1/2x-1 y=1/2x-5

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

You're welcome